A chemist mixes a solution containing 18% alcohol with a second solution containing 45% alcohol. The result is 12 oz of solution that is 36% alcohol. How many ounces of each solution are mixed?
Answer by stanbon(75887)

A chemist mixes a solution containing 18% alcohol with a second solution containing 45% alcohol. The result is 12 oz of solution that is 36% alcohol. How many ounces of each solution are mixed?
-----------------------------
Quantity: x + y = 12 oz
Alcohol :0.18x+0.45y = 0.36*12
------------------------------------
Multiply Quantity by 18
Multiply Alcohol by 100
--------------------------------------
18x + 18y = 18*12
18x + 45y = 36*12
-----
27y = 18*12
----
y = 2*4
y = 8 oz (amt. of 45% solution needed in the mixture)
-------
Since x+y = 12, x = 4 oz. (amt. of 18% needed in the mixture)
========================================================
Cheers,
Stan H.